# Non-equilibrium dynamics of localization phase transition in the non-Hermitian Disorder-Aubry-André model

### v2

@@ Abstract
The driven dynamics of localization transitions in a non-Hermitian Disorder-Aubry-André (DAA) model are thoroughly examined under both open boundary conditions (OBC) and periodic boundary conditions (PBC). Through an analysis of the static properties of observables, including the localization length ($\xi$), inverse participation ratio ($\rm IPR$), and energy gap ($\Delta E$), we found that the critical exponents examined under PBC are also applicable under OBC. The Kibble-Zurek scaling (KZS) for the driven dynamics in the non-Hermitian DAA systems is formulated and numerically verified for different quench directions. Notably, for the dynamical paths considered, boundary conditions had minimal impact on the evolution process. This study generalizes the application of the KZS to the dynamical localization transitions within systems featuring dual localization mechanisms.
@@ I Introduction
The study of localization phase transition behaviors is a longstanding research topic in condensed matter physics Anderson1958 ; Aubry1980 ; Biddle070601 ; Xu2019 ; Khemani2017 ; HepengYao2019 ; Schirmann2024 ; Ribeiro2024 ; Jan2023 ; Alexey2023 ; Agrawal2020 ; Agrawalprl2020 ; Goblot2020 ; Roy2022 ; Agrawal2022 ; Zhou2023 ; Smith176601 . Besides the Anderson model Anderson1958 , the localization phase transition has also been found in the quasi-periodic systems, and various quasi-periodic models have been proposed, such as the Fibonacci model Strkalj2021 ; zhai2021 ; Guo2024 , the Aubry-André (AA) model and its various extensions Aubry1980 ; Sarma1988 ; Biddle2011 ; Wangprl2020 ; Liu2021 ; Gao2024 ; Ting2022 ; Longhi20192 ; Zeng2024 ; Wang2021 ; Borgnia1 ; GYSun2022 . Studies have shown that there are certain differences in the localization phase transition between Anderson model and quasi-periodic models. For example, the localization phase transition can even occur for the one-dimensional (1D) AA model Aubry1980 ; Jan2023 ; Alexey2023 , and the critical exponents for the Anderson and AA models belong to two different universality classes Agrawal2020 ; Wei2019 ; Wang014206 ; Cestari2011 ; Xuan2022 . On the other hand, significant advancements in simulating nonequilibrium dynamics in quantum systems have been made Wangxh2024 ; ChenJJ2024 ; Cheng2399 ; Xue2024 ; Zhang2023131 ; You2024109 ; Mak2024 ; Lizhen2024 . Considerable attention has also been given to the nonequilibrium dynamics in localization transitions Yang2017 ; Heyl2018 ; Bertini2024 ; Khan2024 . Several protocols have been employed to drive a quasi-periodic or disorder system out of equilibrium, such as suddenly changing the Hamiltonian Yang2017 ; Yin2018 ; Modak2021 , introducing a time-(quasi)periodic driving Yang2019 ; Zhou2021 ; ZhouDu2021 , or linearly quenching across the localization transition point Sinha2019 ; Liang2024 ; Xuan2023 . Sudden quenching of the Anderson and AA models, with initial and post-quench Hamiltonians in different phases, will result in periodic zeros in the Loschmidt echo, indicating the localization-delocalization transition Yang2017 ; Yin2018 . The driven dynamics during a linear quench across the localization transition point can be accurately described by the Kibble-Zurek scaling (KZS) Sinha2019 ; Liang2024 ; Xuan2023 .
@@
In recent years, the novel phenomena induced by non-Hermiticity have been extensively studied Longhi2019 ; Luo2021 ; Jing2024 ; Yuto2021 ; Lv2022 ; Jiang2023 ; Han2023 ; Sarkar2023 ; Sunwy2024 ; XiaoL2024 ; Daitx2024 , including the non-Hermitian skin effect (NHSE) Zhang2109431 ; Liu2024 ; Gliozzi2024 ; YoshidaT2024 and exceptional points Wittrock2024 ; Wang024202 ; Martinez2018 ; Aodong706 ; Wingenbach2024 . As discussions on non-Hermitian mechanisms have deepened, researchers have integrated these mechanisms into the study of localization Jiang2019 ; Guo2021 ; Acharya2024 ; Guocx2024 ; Kochergin2024 . The introduction of non-Hermiticity can give rise to a new dimension for tuning localization transitions and uncovers a series of exotic new phenomena zhai2021 ; Jiang2019 ; zhai2020 ; Hatano1996 ; Hatano1998 ; Luo0153523 ; Chen144208 ; Wang024514 ; Longhi224206 ; Jazaeri2001 ; PWang2019 ; Tang2021 ; Chaohua123048 ; YCWang2023 ; KSuthar2022 ; Jiang2024 ; Hamazaki2019 . For instance, the localization transition in the non-Hermitian quasi-periodic system is accompanied by a transition from real to complex eigenvalues, as well as a topological phase transition in the eigenvalue spectrum Liu2021 ; Jiang2019 ; zhai2020 ; Hamazaki2019 . And the NHSE suggests that changes in boundary conditions can result in substantial alterations in the bulk localization properties of non-Hermitian systems Guo2021 . In addition, non-Hermiticity can also lead to critical exponents of the system belonging to different universality classes compared to the Hermitian case Luo2021 ; Zhai2022 .
Figure 1: (a) A schematic representation of the phase diagram for the non-Hermitian DAA model. When $\delta=-2J_{R}$ (denoted by the green point), this model recovers the non-Hermitian Anderson model. The light blue region denotes the localization region, the blue region denotes the critical region, the red point denotes the critical point of the non-Hermitian DAA model. The purple line depicting the skin effect phase under OBC or the extened phase under PBC. The typical spatial distributions of the ground eigenstates for the non-Hermitian DAA model are shown for $\Delta=0$ under (b1) OBC and (b2) PBC. Similarly, the typical spatial distributions of the ground eigenstates for $\Delta=1$ are illustrated under (c1) OBC and (c2) PBC. In these simulations, we use $g=0.5$, $\phi=0$, $\delta=-0.5$, and $L=377$.
@@
More recently, researchers have increasingly focused on exploring the nonequilibrium dynamics of systems with multiple localization mechanisms Liang2024 ; Xuan2023 . Bu et al. introduced a Disorder-Aubry-André(DAA) model, combining disorder and quasi-periodic localization, and investigated its driven dynamics Xuan2022 ; Xuan2023 . Their work proposed a new scaling mechanism, providing a novel perspective on the study of localization phase transition dynamics in systems with coexisting multiple localization mechanisms. However, the interplay between the non-Hermiticity and the non-equilibrium dynamics in such systems remains unexplored. Here, we investigate the driven dynamics of localization transitions in a non-Hermitian DAA model with non-reciprocal hopping under both OBC and PBC. The KZS for the driven dynamics in the non-Hermitian DAA model is constructed and verified numerically across different quench directions. The rest of the paper is arranged as follows: the non-Hermitian DAA model and phase diagram are introduced in Sec. II. The static scaling properties are studied in Sec. III. Then, in Sec. IV, the KZS for the driven dynamics in the non-Hermitian DAA model under OBC and PBC is constructed and numerically verified. A summary is given in Sec. V.
@@ II The non-Hermitian DAA Model and phase diagram
Here, $c_{j}^{\dagger}(c_{j})$ are creation (annihilation) operator of the hard-core boson and $\gamma$ is an irrational number. $J_{L}=Je^{-g}$ and $J_{R}=Je^{g}$ are the asymmetry hopping coefficients. $\Delta$ and $(2J_{R}+\delta)$ measure the the amplitude of disorder and the quasi-periodic potential, respectively. $w_{j}\in[-1,1]$ gives the quenched disorder configuration, $\phi\in[0,1)$ is phase of the potential. We assume $J=1$ as the unity of energy, and $\gamma=(\sqrt{5}-1)/2$. To satisfy PBC of the quasi-periodic potential, we approximate $\gamma$ as a rational number $F_{n}/F_{n+1}$ where $F_{n+1}=L$ and $F_{n}$ are the Fibonacci numbers Jiang2019 ; Zhai2022 .
A schematic phase diagram of the non-Hermitian DAA model near the critical point is illustrated in Fig. 1(a). For $\delta=-2J_{R}$, this model returns to the non-Hermitian Anderson model Aubry1980 . However, when $\delta=0$, any finite $\Delta$ can also cause localization at the critical point of the non-Hermitian AA model, but the localization characteristics in this case differ from those of the purely non-Hermitian Anderson model. Therefore, for the non-Hermitian DAA model, the point $(\delta,\Delta)=(0,0)$ represents its critical point, and the critical region surrounding this point is jointly defined by $\Delta$ and $\delta$.
@@
When $\Delta=0$, this model recovers the non-Hermitian AA model Jiang2019 . For $\delta>0$, all the eigenstates are localized under PBC and OBC. For $\delta<0$, all the eigenstates are extended under PBC Zhai2022 , whereas it exhibits the skin-effect phase under OBC Jiang2019 ; Zhaifphy . As illustrated in Figs. 1(b1) and (b2), the spatial distribution of the right eigenvector of the ground states ($|\Psi(j)\rangle$) with $\delta<0$ and $\Delta=0$ under OBC and PBC are depicted, respectively. Here, the ground state is defined as the eigenstate possessing the smallest real eigenenergy among all possible states. It is shown that the wave function is localized on the boundary under OBC, but it is evenly distributed within a small range under PBC. For $\delta>0$, the system will be in a localized state, regardless of the value of $\Delta$. For $\delta<0$, any finite $\Delta$ can also cause the system to transition into a localized state. As illustrated in Figs. 1(c1) and (c2), when $\Delta=1$ and $\delta<0$, the spatial distribution of localized states exhibits a high degree of similarity under both boundary conditions.
To discuss the impact of disorder on the localization phase transition in the non-Hermitian AA model, we further explore the energy spectrum of the non-Hermitian DAA model. Under PBC, the energy spectrum for the non-Hermitian AA model under PBC does not undergo significant changes when disorder is introduced, but it disrupts the correspondence between the transition from real to complex of energy spectrum and the localization transition Sun2024 . Here, Fig. 2 displays the energy spectrum of the non-Hermitian DAA model under OBC. The results show that, unlike the fully real energy spectrum of the non-Hermitian AA model under OBC Jiang2019 , the introduction of disorder leads to the emergence of imaginary parts in the energy spectrum in the region where $\delta<0$.
@@ III Static critical properties of the Non-Hermitian DAA model
III.1 Static scaling forms in the critical region
In this section, we study the static properties within the critical region, utilizing the localization length ($\xi$), the inverse participation ratio ($\rm IPR$), and the energy gap between the ground state and the first excited state ($\Delta E$).
@@ III.1 Static scaling forms in the critical region
$\xi=\sqrt{\sum_{n>n_{c}}^{L}[(n-n_{c})^{2}]P_{i}},$ (2)
where $P_{i}$ denotes the probability of the wavefunction at site $i$, and $n_{c}\equiv\sum nP_{i}$ represents the localization center. In the vicinity of a critical point, $\xi$ scales with the distance to the critical point $\varepsilon$ as Aubry1980
$\xi\propto\varepsilon^{-\nu},$ (3)
where $\nu$ is a critical exponent. For the pure non-Hermitian AA model, under both PBC and OBC, $\varepsilon=\delta$ and $\nu=\nu_{\delta}=1$ Zhai2022 ; Zhaifphy . For the pure non-Hermitian Anderson model, $\varepsilon=\Delta$ and $\nu=\nu_{A}=1.99$ Sun2024 . For the non-Hermitian DAA model, $\Delta$ constitutes another significant factor at $(\delta,\Delta)=(0,0)$ Xuan2022 ; Sun2024 . Specifically, when $\delta=0$ in Eq. (II), $\varepsilon=\Delta$ and $\nu=\nu_{\Delta}=0.52$ under PBC Sun2024 . $\nu_{\Delta}$ is a new parameter distinct from that of the non-Hermitian Anderson model and the non-Hermitian AA model.
@@
In the extended states, the scaling of ${\rm IPR}$ is given by ${\rm IPR}\propto L^{-1}$. For the localized states or skin effect sates, the scaling is ${\rm IPR}\propto L^{0}$ Jiang2019 . At the critical point, the $\rm IPR$ scales as
${\rm IPR}\propto L^{-s/\nu},$ (5)
where $s$ is a critical exponent. As $L$ approaches infinity, the ${\rm IPR}$ scales with the distance to the critical point $\varepsilon$ according to
${\rm IPR}\propto\varepsilon^{s}.$ (6)
For the non-Hermitian AA model, under both OBC and PBC, it is found that $\varepsilon=\delta$ and $s=s_{\delta}=0.1197$ Zhai2022 ; zhai2021 .
At the critical point of localization transition, the energy gap $\Delta E$ exhibits a scaling relationship with the lattice size $L$, given by
$\Delta E\propto L^{-z},$ (7)
where $z$ is a critical exponent. For the non-Hermitian AA model, under both OBC and PBC, it is found that $\varepsilon=\delta$ and $z=z_{\delta}=2$ Zhai2022 . when $L$ tends to infinity, energy gap $\Delta E$ scales as
$\Delta E\propto\varepsilon^{\nu z}.$ (8)
Considering the finite-size effect, the general scaling formulation for a quantity $Y$ referring to a single parameter can be expressed as:
$Y(\varepsilon)$ $=$ $L^{y/\nu}f(\varepsilon L^{1/\nu}),$ (9)
@@
where $y$ is the critical exponent of $Y$, defined such that $Y\propto\varepsilon^{-y}$ as $L\rightarrow\infty$. The function $f(.)$ represents the scaling function. By taking $\Delta$ and $\delta$ as scaling variables simultaneously, the comprehensive finite-size scaling expressions for the three observables are given by
$\xi$ $=$ $Lf_{1}(\delta L^{1/\nu_{\delta}},\Delta L^{1/\nu_{\Delta}}),$ (10)
${\rm IPR}$ $=$ $L^{-s_{\delta}/\nu_{\delta}}f_{2}(\delta L^{1/\nu_{\delta}},\Delta L^{1/\nu_{\Delta}}),$ (11)
${\Delta E}$ $=$ $L^{-z_{\delta}}f_{3}(\delta L^{1/\nu_{\delta}},\Delta L^{1/\nu_{\Delta}}).$ (12)
These scaling functions, Eqs. (10) to (12), are applicable within the critical region of the non-Hermitian DAA model.
Figure 3: Scaling properties in the ground state for fixed $\delta L^{1/\nu_{\delta}}=1$ under OBC. The curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaled for different $L^{\prime}$s. The curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaled for different $L^{\prime}$s. The curves of $\Delta E$ versus $\Delta$ before (c1) and after (c2) rescaled for different $L^{\prime}$s. Here, $g=0.5$, and the result is averaged for $1000$ samples. The double-logarithmic scales are used.
III.2 Numerical results of static scaling of the non-Hermitian DAA model
@@
The critical exponents of the non-Hermitian DAA model under PBC have been examined Sun2024 . It is found that the critical exponents vary along the $\delta$ and $\Delta$ directions because these parameters represent two distinct relevant directions. Under PBC, the critical exponents along $\delta$ direction with $\Delta=0$ are $(\nu,s,z)=(\nu_{\delta},s_{\delta},z_{\delta})=(1,0.1197,2)$ Zhai2022 . In contrast, the scaling exponents along the $\Delta$ direction are $(\nu,s,z)=(\nu_{\Delta},s_{\Delta},z_{\Delta})=(0.52,0.0642,2)$.
@@
Figure 4: Scaling properties in the ground state for fixed $\delta L^{1/\nu_{\delta}}=-1$ under OBC. The curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaled for different $L^{\prime}$s. The curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaled for different $L^{\prime}$s. The curves of $\Delta E$ versus $\Delta$ before (c1) and after (c2) rescaled for different $L^{\prime}$s. Here, $g=0.5$, and the result is averaged for $1000$ samples. The double-logarithmic scales are used.
@@
Here, we study the static scaling properties of the non-Hermitian DAA model under OBC by employing the exponents obtained from fitting under PBC. We numerically validate Eqs. (10) to (12) under OBC by maintaining $\delta L^{1/\nu_{\delta}}$ at a constant value. Fig. 3 depict the scaling behavior of $\xi$, $\rm IPR$ and $\Delta E$ as functions of $\Delta$ for $\delta L^{1/\nu_{\delta}}=1$. After rescaling these values according to Eqs. (10) to (12), the resulting curves align closely, thereby confirming the validity of these equations. For the critical region where $\delta<0$, similar numerical results are presented in Fig. 4 for $\delta L^{1/\nu_{\delta}}=-1$. The alignment of the rescaled curves shown in Figs. 4 (a2), (b2) and (c2) further verifies Eqs. (10) to (12).
These results confirm that the same set critical exponent of the non-Hermitian DAA model under PBC is typically sufficient to characterize the critical behavior in the critical region under OBC.

### v3

# Non-equilibrium dynamics of localization phase transition in the non-Hermitian Disordered Aubry-André model

@@ Abstract
The driven dynamics of localization transitions in a non-Hermitian Disordered Aubry-André (DAA) model are examined under both open boundary conditions (OBC) and periodic boundary conditions (PBC). Through an analysis of the static properties of observables, including the localization length ($\xi$), inverse participation ratio ($\rm IPR$), and energy gap ($\Delta E$), we found that the critical exponents examined under PBC are also applicable under OBC. The Kibble-Zurek scaling (KZS) for the driven dynamics in the non-Hermitian DAA systems is formulated and numerically verified for different local-to-local quench directions. The hybrid KZS (HKZS) in the overlapping critical region of non-Hermitian DAA and Anderson localization is proposed and numerically confirmed the validity across a local-to-skin quench direction. This study generalizes the application of the KZS to the dynamical localization transitions within systems featuring dual localization mechanisms.
@@ I Introduction
The study of localization phase transition behaviors is a longstanding research topic in condensed matter physics Anderson1958 ; Aubry1980 ; Biddle070601 ; Xu2019 ; Khemani2017 ; HepengYao2019 ; Schirmann2024 ; Ribeiro2024 ; Jan2023 ; Alexey2023 ; Agrawal2020 ; Agrawalprl2020 ; Goblot2020 ; Roy2022 ; Agrawal2022 ; Zhou2023 ; Smith176601 . Besides the Anderson model Anderson1958 , the localization phase transition has also been found in the quasi-periodic systems, and various quasi-periodic models have been proposed, such as the Fibonacci model Strkalj2021 ; zhai2021 ; Guo2024 , the Aubry-André (AA) model and its various extensions Aubry1980 ; Sarma1988 ; Biddle2011 ; Wangprl2020 ; Liu2021 ; Gao2024 ; Ting2022 ; Longhi20192 ; Zeng2024 ; Wang2021 ; Borgnia1 ; GYSun2022 . Studies have shown that there are certain differences in the localization phase transition between Anderson model and quasi-periodic models. For example, the localization phase transition can even occur for the one-dimensional (1D) AA model Aubry1980 ; Jan2023 ; Alexey2023 , and the critical exponents for the Anderson and AA models belong to two different universality classes Agrawal2020 ; Wei2019 ; Wang014206 ; Cestari2011 ; Xuan2022 . These two models have been realized and investigated on various experimental platforms Billy2008 ; Linrj2023 ; xueprl2022 ; xueNC2022 , including RLC circuit lattices, acoustic and photonic lattices, single-photon quantum walks, mechanical metamaterials, and cold atoms.
On the other hand, significant advancements in simulating nonequilibrium dynamics in quantum systems have been made Wangxh2024 ; ChenJJ2024 ; Cheng2399 ; Xue2024 ; Zhang2023131 ; You2024109 ; Mak2024 ; Lizhen2024 . Considerable attention has also been given to the nonequilibrium dynamics in localization transitions Yang2017 ; Heyl2018 ; Bertini2024 ; Khan2024 . Several protocols have been employed to drive a quasi-periodic or disorder system out of equilibrium, such as suddenly changing the Hamiltonian Yang2017 ; Yin2018 ; Modak2021 , introducing a time-(quasi)periodic driving Yang2019 ; Zhou2021 ; ZhouDu2021 , or linearly quenching across the localization transition point Sinha2019 ; Liang2024 ; Xuan2023 . Sudden quenching of the Anderson and AA models, with initial and post-quench Hamiltonians in different phases, will result in periodic zeros in the Loschmidt echo, indicating the localization-delocalization transition Yang2017 ; Yin2018 .The driven dynamics during a linear quench across the localization transition point can be accurately described by the Kibble-Zurek scaling (KZS) Sinha2019 ; Liang2024 ; Xuan2023 .
@@
In recent years, the novel phenomena induced by non-Hermiticity have been extensively studied Longhi2019 ; Luo2021 ; Jing2024 ; Yuto2021 ; Lv2022 ; Jiang2023 ; Han2023 ; Sarkar2023 ; Sunwy2024 ; XiaoL2024 ; Daitx2024 , including the non-Hermitian skin effect (NHSE) Zhang2109431 ; Liu2024 ; Gliozzi2024 ; YoshidaT2024 and exceptional points Wittrock2024 ; Wang024202 ; Martinez2018 ; Aodong706 ; Wingenbach2024 . As discussions on non-Hermitian mechanisms have deepened, researchers have integrated these mechanisms into the study of localization Jiang2019 ; Guo2021 ; Acharya2024 ; Guocx2024 ; Kochergin2024 . The introduction of non-Hermiticity can give rise to a new dimension for tuning localization transitions and uncovers a series of exotic new phenomena zhai2021 ; Jiang2019 ; zhai2020 ; Hatano1996 ; Hatano1998 ; Luo0153523 ; Chen144208 ; Wang024514 ; Longhi224206 ; Jazaeri2001 ; PWang2019 ; Tang2021 ; Chaohua123048 ; YCWang2023 ; KSuthar2022 ; Jiang2024 ; Hamazaki2019 . For instance, the localization transition in the non-Hermitian quasi-periodic system is accompanied by a transition from real-to-complex eigenvalues, as well as a topological phase transition in the eigenvalue spectrum Liu2021 ; Jiang2019 ; zhai2020 ; Hamazaki2019 . And the NHSE suggests that changes in boundary conditions can result in substantial alterations in the bulk localization properties of non-Hermitian systems Guo2021 . In addition, non-Hermiticity can also lead to critical exponents of the system belonging to different universality classes compared to the Hermitian case Luo2021 ; Zhai2022 . Experimentally, the introduction of non-Hermiticity can be achieved by using cold atomic traps Linrj2023 , specific optical elementsxueprl2022 ; xueNC2022 , or microwave circuit components Jiang2019 . For example, in the single-photon interference network experimental platform, one can simulate different quantum states by encoding the polarization state of photons, and control the selective loss of the polarization state of photons on different paths by setting the angle of the half-wave plate, which reduces the probability amplitude of the system on certain paths and thus induces non-Hermiticity xueprl2022 ; xueNC2022 .
@@
More recently, researchers have increasingly focused on exploring the nonequilibrium dynamics of systems with multiple localization mechanisms Liang2024 ; Xuan2023 . Bu et al. introduced a Disordered Aubry-André(DAA) model, combining disorder and quasi-periodic localization, and investigated its driven dynamics Xuan2022 ; Xuan2023 . Their work proposed a new scaling mechanism, providing a novel perspective on the study of localization phase transition dynamics in systems with coexisting multiple localization mechanisms. However, the interplay between the non-Hermiticity and the non-equilibrium dynamics in such systems remains unexplored. Here, we investigate the driven dynamics of localization transitions in a non-Hermitian DAA model with non-reciprocal hopping under both OBC and PBC. The KZS for the driven dynamics in the non-Hermitian DAA model is constructed and verified numerically across different quench directions. The rest of the paper is arranged as follows: the non-Hermitian DAA model and phase diagram are introduced in Sec. II. The static scaling properties are studied in Sec. III. In Sec. IV, the KZS for the driven dynamics in the non-Hermitian DAA model under OBC and PBC is constructed and numerically verified. Then, in Sec. V, the hybrid KZS (HKZS) in the overlapping critical region of non-Hermitian DAA and Anderson localization is proposed and numerically confirmed the validity across a local-to-skin quench direction. Sec. VI introduces an experimental scheme for implementing the non-Hermitian DAA model through time-multiplexed photonic quantum walk system. A summary is given in Sec. VII.
@@ II The non-Hermitian DAA Model and phase diagram
Here, $c_{j}^{\dagger}(c_{j})$ are creation (annihilation) operator of the hard-core boson and $\gamma$ is an irrational number. $J_{L}=Je^{-g}$ and $J_{R}=Je^{g}$ are the asymmetry hopping coefficients. The parameter $g$ characterizes the extent of asymmetric hopping between lattice sites within the system. A larger $g$ value corresponds to a greater disparity in asymmetric hopping between the left and right directions. The adjustment of the value of $g$ can also bring about localized phase transitions and change the degree of localization zhai2021 ; Zhaifphy . Notably, $g$ plays a crucial role in altering the critical point separating the extended phases from the localization phase Jiang2019 . When studying the scaling invariance near the critical point of a non-Hermitian system and determining the relevant critical exponents, it is necessary to first determine the value of $g$ in order to locate the phase transition point. However, previous research has demonstrated that, within a specific range of values, $g$ does not significantly influence the magnitude of the critical exponents Zhai2022 . Consequently, in this work, we treat $g$ as a fixed parameter for introducing non-Hermiticity, with a predetermined value of 0.5. $\Delta$ and $(2J_{R}+\delta)$ measure the amplitude of disorder and the quasi-periodic potential, respectively. $w_{j}\in[-1,1]$ gives the quenched disorder configuration, $\phi\in[0,1)$ is phase of the potential. We assume $J=1$ as the unity of energy, and $\gamma=(\sqrt{5}-1)/2$. To satisfy PBC of the quasi-periodic potential, we approximate $\gamma$ as a rational number $F_{n}/F_{n+1}$ where $F_{n+1}=L$ and $F_{n}$ are the Fibonacci numbers Jiang2019 ; Zhai2022 .
Figure 1: (a) A schematic representation of the phase diagram for the non-Hermitian DAA model. When $\delta=-2J_{R}$ (denoted by the green point), this model recovers the non-Hermitian Anderson model. The light blue region denotes the localization region of non-Hermitian DAA model, the blue region denotes the critical region of non-Hermitian DAA model, the red point denotes the critical point of the non-Hermitian DAA model. The purple region denotes the critical region of non-Hermitian Anderson model. The part where the blue region and the purple region overlap is the overlapping critical region, in which the critical regions of non-Hermitian DAA and non-Hermitian Anderson localization coexist. When $\delta<0$ and $\Delta=0$, all the eigenstates are extended under PBC, whereas it exhibits the skin-effect phase under OBC. The typical spatial distributions of the state with the lowest real part of the eigenenergy for the non-Hermitian DAA model are shown for $\Delta=0$ under (b1) OBC and (b2) PBC. Similarly, the typical spatial distributions of the state with the lowest real part of the eigenenergy for $\Delta=1$ are illustrated under (c1) OBC and (c2) PBC. In these simulations, we use $g=0.5$, $\phi=0$, $\delta=-0.5$, and $L=377$.
@@
A schematic phase diagram of the non-Hermitian DAA model near the critical point is illustrated in Fig. 1(a). When $\delta>0$, the system will be in a localized phase regardless of the value of $\Delta$. When $\delta<0$ and $\Delta=0$, all the eigenstates are extended under PBC Zhai2022 , whereas it exhibits the skin-effect phase under OBC Jiang2019 ; Zhaifphy . As illustrated in Figs. 1(b1) and (b2), the spatial distribution of the right eigenvector of the state with the lowest real part of the eigenenergy ($|\Psi(j)\rangle$) with $\delta<0$ and $\Delta=0$ under OBC and PBC are depicted, respectively. It is shown that the wave function is localized on the boundary under OBC, but it is evenly distributed within a small range under PBC. When $\delta<0$, any non-zero value of $\Delta$ can induce localization under PBC, whereas a competitive interplay emerges between the skin-effect state and the localized state under OBC. As shown in Figs. 1(c1) and (c2), when $\Delta=1$ and $\delta<0$, the spatial distributions of states exhibit highly similar localization characteristics under both boundary conditions.
The point $(\delta,\Delta)=(0,0)$ represents the critical point of the non-Hermitian DAA model, and the critical region surrounding this point is jointly defined by $\Delta$ and $\delta$. In this critical region, the system is also in a localized phase, and the correlation length, which characterizes the degree of localization, will exhibit an exponential divergence trend as the parameter $\delta$ or $\Delta$ changes. Moreover, for $\delta<0$ and infinitesimal $\Delta$, there is a critical region of Anderson localization. Therefore, for $\delta<0$, the critical region of non-Hermitian Anderson localization and the critical region of non-Hermitian DAA will inevitably overlap with each other, thereby forming an overlapping region.
@@
To discuss the impact of disorder on the localization phase transition in the non-Hermitian AA model, we further explore the energy spectrum of the non-Hermitian DAA model. The value of $g$ and the magnitude of the potential together lead to the appearance of complex energy spectra in the extended state region of the system’s energy spectrum. Under PBC, the energy spectrum for the non-Hermitian AA model under PBC does not undergo significant changes when disorder is introduced, but it disrupts the correspondence between the transition from real to complex of energy spectrum and the localization transition Sun2024 . Here, Fig. 2 displays the energy spectrum of the non-Hermitian DAA model under OBC. The results show that, unlike the fully real energy spectrum of the non-Hermitian AA model under OBC Jiang2019 , the introduction of disorder leads to the emergence of imaginary parts in the energy spectrum in the region where $\delta<0$.
@@ III Static critical properties of the Non-Hermitian DAA model
In this section, we study the static properties within the critical region, utilizing the localization length ($\xi$), the inverse participation ratio ($\rm IPR$), and the energy gap ($\Delta E$) between the lowest-real-eigenenergy state and the second-lowest-real-eigenenergy state. $\Delta E$ is a key quantity for understanding the stability, dynamics, phase transitions, and potential applications of non-Hermitian systems.
@@
$\xi=\sqrt{\sum_{j>j_{c}}^{L}[(j-j_{c})^{2}]P_{j}},$ (2)
where $P_{j}=||\Psi(j)\rangle|^{2}$ denotes the probability of the wavefunction at site $j$, and $j_{c}\equiv\sum jP_{j}$ represents the localization center.
@@
In the extended states, the scaling of $\rm IPR$} is given by ${\rm IPR}\propto L^{-1}$. For the localized states or skin effect sates, the scaling is ${\rm IPR}\propto L^{0}$ Jiang2019 .
@@
By taking $\Delta$ and $\delta$ as scaling variables simultaneously, the comprehensive finite-size scaling expressions for the three observables are given by
$\xi$ $=$ $Lf_{1}(\delta L^{\frac{1}{\nu_{\delta}}},\Delta L^{\frac{1}{\nu_{\Delta}}}),$ (4)
${\rm IPR}$ $=$ $L^{-\frac{s_{\delta}}{\nu_{\delta}}}f_{2}(\delta L^{\frac{1}{\nu_{\delta}}},\Delta L^{\frac{1}{\nu_{\Delta}}}),$ (5)
${\Delta E}$ $=$ $L^{-z_{\delta}}f_{3}(\delta L^{\frac{1}{\nu_{\delta}}},\Delta L^{\frac{1}{\nu_{\Delta}}}).$ (6)
@@
The function $f_{i}(.)$ represents the scaling function. These scaling functions, Eqs. (4) to (6), are applicable within the critical region of the non-Hermitian DAA model. Research has found that these two parameters represent two distinct relevant directions, and the critical exponents vary along the $\delta$ or $\Delta$ directions Xuan2023 ; Xuan2022 ; Sun2024 . The critical exponents along $\delta$ direction with $\Delta=0$ are $(\nu_{\delta},s_{\delta},z_{\delta})=(1,0.1197,2)$, under both PBC and OBC Zhai2022 ; zhai2021 ; Zhaifphy . In contrast, the scaling exponents along the $\Delta$ direction are $(\nu_{\Delta},s_{\Delta},z_{\Delta})=(0.52,0.0642,2)$ under PBC Sun2024 .
@@
Figure 3: Scaling properties in the state with the lowest real part of the eigenenergy for fixed $\delta L^{1/\nu_{\delta}}=-1$ under OBC. The curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaled for different $L^{\prime}$s. The curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaled for different $L^{\prime}$s. The curves of $\Delta E$ versus $\Delta$ before (c1) and after (c2) rescaled for different $L^{\prime}$s. Here, $g=0.5$, and the result is averaged for 10001000 samples. The double-logarithmic scales are used.
@@
Here, we first verify the applicability of the exponents obtained from fitting along the $\Delta$ direction under PBC to the non-Hermitian DAA model under OBC. We numerically validate Eqs. (4) to (6) under OBC by maintaining $\delta L^{1/\nu_{\delta}}$ at a constant value. Fig. 3 depict the scaling behavior of $\xi$, $\rm IPR$ and $\Delta E$ as functions of $\Delta$ for $\delta L^{1/\nu_{\delta}}=-1$. After rescaling these values according to Eqs. (4) to (6), the resulting curves align closely, thereby confirming the validity of these equations. Setting $\delta L^{1/\nu_{\delta}}$ to -1 is an arbitrary assignment. The simulation results for another randomly selected value of -0.5, which also yielded similar outcomes, are presented in Appendix A. These results confirm that the same set critical exponent of the non-Hermitian DAA model under PBC is typically sufficient to characterize the critical behavior in the critical region under OBC.